### v2

@@ Abstract
Future experiments at hadron colliders require an evolution of the tracking sensors to ensure sufficient radiation hardness as well as space and time resolution to handle unprecedented particle fluxes. 3D diamond sensors with laser-graphitized electrodes are promising candidates due to their strong binding energy, small atomic number, and high carrier mobility. However, the high resistance of the engraved electrodes delays the propagation of the induced signals towards the readout electronics, thereby degrading the precision of the timing measurements. So far, this effect has been the dominant factor limiting the time resolution of these devices, with other contributions, such as those due to electric field inhomogeneities or electronic noise, typically neglected. Recent advancements in graphitization technology, however, motivate a renewed effort in modeling signal generation in 3D diamond detectors, to achieve more reliable predictions. To this purpose, we apply an extended version of the Ramo-Shockley theorem describing the effect of signal propagation as a time-dependent weighting potential, obtained by solving numerically the Maxwell’s equations in a quasi-static approximation. We developed a custom spectral method solver and validated it against COMSOL MultiPhysics®. The response of the modeled sensor to a beam of particles is then simulated using Garfield++ and is compared to the data acquired in a beam test carried on in 2021 by the TimeSPOT Collaboration at the SPS, at CERN. Based on the results obtained with this simulation workflow, we conclude that reducing the resistivity of the graphitic columns remains the priority for significantly improving the time resolution of 3D diamond detectors. Once achieved, optimization of the detector geometry and readout electronics design will become equally important steps to further enhance the timing performance of these devices.
@@ 1 Introduction
The development of the ATLAS and CMS detectors for the High-Luminosity Large Hadron Collider [1, 2, 3, 4], the upcoming upgrade of the LHCb experiment [5, 6], and the detectors at the hadronic version of the Future Circular Collider (FCC-hh) [7], or at the Muon Collider [8] exemplify the growing demand for sensors capable of withstanding an unprecedented flux of ionizing particles.
@@
Similar challenges arise in technological and applied research, such as in the development of dosimeters for Flash Radiotherapy [9, 10] or in providing precise process monitors in nuclear fission and fusion facilities [11].
@@
Physics requirements for experiments at future high luminosity colliders impose severe constraints especially on tracking detectors placed close to the interaction points. Besides very high radiation hardness they must feature excellent time and space resolution to ensure an efficient tracking and high precision in primary and secondary vertex reconstruction [6].
@@
Among the many technologies that are being considered, like for example the recently proposed doping compensation in LGAD silicon sensors [12], diamond detectors have been extensively studied as highly promising candidates for addressing the challenges posed by extreme operating conditions. Diamond sensors offer indeed several advantages. Firstly, they have enhanced radiation tolerance due to the strong lattice binding energy and relatively low atomic number. Secondly, the high mobility and saturation velocity of charge carriers imply faster signal collection and improved time resolution compared to silicon devices [13, 14]. Moreover, diamond sensors have a sufficiently large band-gap to make them solar-blind and operable at higher temperatures. This characteristic is particularly advantageous for applications where cooling systems are impractical.
@@
An intriguing feature of diamond is its ability to undergo laser-induced graphitization, transforming few of diamond around the laser focus into a conductive admixture of graphite and amorphous carbon. This process enables the fabrication of full-carbon sensors with high flexibility in the geometry of electrodes which can be realized either on the surface or in the bulk of the diamond crystals [15, 16, 17, 18].
@@
Similarly to the silicon sensors, the 3D geometry, where electrodes are perpendicular to the sensor surface, offers significant advantages over the traditional planar geometry. In this configuration the drift length of charge carriers is substantially shorter with respect to the planar geometry reducing the probability of charge trapping by radiation induced defects and shortening the charge collection time leading to faster signals. A sketch of a 3D electrode arrangement and a picture of a 3D diamond sensor are shown in Figure 2 and Figure 2 respectively.
@@
The fabrication of 3D diamond sensors presents substantial challenges due to the high refractive index of diamond, which degrades the laser focus due to spherical aberration, and because of the major resistance of the diamond bulk to laser-induced phase transitions. The transition from diamond to graphite within the bulk is complicated by differences in density between the two materials, leading to mechanical strain in crystal lattice. Despite these challenges, significant progress has been made. For instance, Salter et al. demonstrated in 2017 that phase transitions in the diamond bulk occur through discrete nano-scale structural modifications aligned with the diagonals of the diamond lattice [19]. The electrical resistivity of these graphitized regions is influenced by random connections between conductive structures and is typically orders of magnitude higher than that of graphitized surfaces [20].
@@
Several other studies, across various scientific communities, proposed methods to reduce the resistivity of the engraved structures. The most promising results have been achieved using femtosecond infrared lasers, where spherical aberration introduced by the diamond surface is corrected through Spatial Light Modulators [21, 22]. This approach has enabled the fabrication of structures parallel to the diamond surface with a recorded resistivity of 0.02 [23]. Additionally, the same experimental setup has demonstrated the capability to engrave multiple electrodes simultaneously [24]. On the other hand, employing Bessel femtosecond-pulsed beams without adaptive spherical aberration correction has resulted in the formation of fine electrodes orthogonal to the diamond surface, exhibiting a resistivity of 0.04 [25].
@@
High resistivity in graphitic structures leads to considerable electrode resistance in diamond sensors. The increased resistance affects the signal propagation time through the electrodes, which becomes the dominant limiting effect of the sensor time resolution. This effect allowed so far to neglect all the contributions to the timing performance due to charge carrier propagation through the bulk as well as those coming from electronic noise. Recently, however, 3D laser-graphitized diamond sensors with a time resolution significantly below 100 ps have been reported [21], strongly motivating a renewed effort to model the signal formation in diamond detectors by properly taking into account the combined effect of carrier transport through the bulk and signal propagation through resistive electrodes.
@@
In this work, we propose an innovative numerical model based on an extension of the Ramo-Shockley theorem to conductive media to simulate the signal induction process in 3D diamond sensors with graphitic electrodes, including the contribution of the resistive elements as described in Refs. [26, 27, 28, 29, 30].
To validate our proposed method, we use experimental data obtained from a beam test conducted at CERN, thereby ensuring the reliability and applicability of our model.
@@
Having validated the method, we simulated sensors with different geometries and configurations to study their time resolution. The simulations allowed us to disentangle the various contributions to the time resolution, enabling a systematic investigation aimed at identifying potential optimizations.
@@ 2 Signal simulation in 3D diamond sensors with graphitic electrodes
Particles traversing a radiation detector release part of their energy inside the detection medium generating charge carriers which drift towards the electrodes inducing a current signal. Simulating the generation of these signals includes therefore three main phases: the interaction of a ionizing particle with the detector medium, the transport of generated charge carriers through the medium and the modeling of the induced signal. If the readout electrodes are significantly resistive, the propagation of the induced signal through the electrodes has to be also properly taken into account. In our simulation workflow we simulate the particle interaction with the detector and the charge carrier transport using the Garfield++ software package [31], using an extension of the Ramo-Shockley theorem to describe the signal generation, taking into account the resistivity of the graphitic electrodes.
The sensor considered for the simulation is a 3D diamond sensor, 500 thick, with 12 -diameter graphitized electrodes, organized on a two-dimensional, -pitch pixel matrix. The readout electrodes are arranged in square cells with polarization electrodes placed at the four corners and the readout electrode at the center. Both polarization and readout electrodes are 450 long and are engraved starting from opposite sides of the crystal. Their resistance was indirectly measured to be 30 k, corresponding to a resistivity of 0.75 . Its response was investigated through simulations of tracks produced by 180 GeV/ pions impinging on the sensor surface.
@@ 2.1 Computing trajectories of charge carriers with Garfield++
The numerical model proposed in this work is built upon established methodologies for simulating the migration of charge carriers generated by the passage of ionizing particles. The motion of charges occurs within the electrostatic field () defined by the polarization voltage applied () to sensor electrodes.
The ionization charge generated by the traversing particle is simulated using the High Energy Electrodynamics (Heed) toolkit [32, 33] as integrated in Garfield++. The drift of generated electrons and holes is determined by the electrostatic field, which determines both their trajectories toward the electrodes and their velocities, calculated by accounting for the carrier mobility () and saturation velocity () in diamond:
Due to the complex structure of the graphitized electrodes, carrier drift is neglected within the regions representing the conductive carbon phase. It should be noted that this simplification introduces discrepancies when the ionizing particle traverses an electrode because the deposited charge is not accounted for in our simulation model. However, beam test results on 3D diamond sensors show tracking efficiencies above 99%, indicating that the charge deposited in the graphitized material effectively contributes to the signals collected at the electrode ends [21].
@@
The results described below were obtained computing the electrostatic potential in the diamond sensor with either a Finite Element Method implemented in COMSOL MultiPhysics® [34] on a custom mesh [30] or using an iterative spectral method on a uniform voxel grid, resetting the boundary conditions on voltage and charge density at each iteration.
@@
2.2 Approximating the effect of resistive electrodes with a discrete impedance network
As a preliminary step, in order to have a simplified benchmark available to gauge our results with, the read out graphitic electrode is described as a discrete impedance network, where the current signal induced by the drift of charge carriers inside the sensor is injected and propagated to the readout electronics input. To calculate the induced current the classical Ramo-Shockley formalism is used [35, 36].
According to the Ramo-Shockley theorem, the current induced on a given electrode by the drift of a charge along a path with velocity , defined in Equation 2.1, is given by
Here is the weighting field, obtained by solving Laplace’s equation for the sensor, where the charge is removed, a fixed potential is applied to the electrode of interest, and all other electrodes are set to ground. The weighting field is time-independent and will be referred to as the steady weighting field in the following.
@@ 2.2 Approximating the effect of resistive electrodes with a discrete impedance network
As already introduced, however, the graphitic electrodes have a sizable resistance and the time constant of the system composed of two oppositely charged graphite electrodes embedded in a diamond dielectric is of the order of 100 ps or larger. A crude approximation to take into account the effect of signal propagation through the graphitized electrodes is to model them as a discrete ladder of capacitors and resistors, as shown in Figure 5.
@@
From the equivalent circuit shown in Figure 5, a delta response function is obtained by applying an elementary pulse in the current generators and measuring the current flowing towards ground. To account for the fluctuations of the ionization charge deposit through the depth of the columns, the elementary current pulse is split randomly through the discrete elements of the model according to the Landau distribution [37]. The transfer function is computed using the Ngspice software package [38] and is then convolved numerically with the signal obtained with Garfield++ using the steady weighting field.
This simple model has several limitations but can still provide useful information on how the signal shape depends on electrode resistance. First of all one has to consider the large uncertainties on the capacitance and the approximations coming in particular from neglecting the correlation effects between charge-carrier trajectories and signal propagation. In addition, the model is valid only insofar as the bias and readout electrodes can be considered symmetric and that the shaping time constant of the readout electronics is of the same order of magnitude of the charge-carrier transport timescale, or larger. These constraints make this model applicable only as a semi-quantitative benchmark while to make systematic studies on detector geometry and configuration optimization a more detailed method shall be used.
Hence, to properly model the signal formation in the diamond sensors with graphitized electrodes we pursue a more general and rigorous approach that in the simple cases described above is expected to yield results consistent with the discrete impedance network model, but applicable to a much wider set of detector configurations and geometries.

### v3

@@ Abstract
Future experiments at hadron colliders require an evolution of the tracking sensors to ensure sufficient radiation hardness as well as space and time resolution to handle unprecedented particle fluxes. 3D diamond sensors with laser-graphitized electrodes are promising candidates due to their strong binding energy, small atomic number, and high carrier mobility. However, the high resistance of the engraved electrodes delays the propagation of the induced signals towards the readout electronics, thereby degrading the precision of the timing measurements. So far, this effect has been the dominant factor limiting the time resolution of these devices, with other contributions, such as those due to electric field inhomogeneities or electronic noise, typically neglected. Recent advancements in graphitization technology, however, motivate a renewed effort in modeling signal generation in 3D diamond detectors, to achieve more reliable predictions. To this purpose, we apply an extended version of the Ramo-Shockley theorem, describing the effect of signal propagation as a time-dependent weighting potential, obtained by numerically solving the Maxwell’s equations in a quasi-static approximation. We developed a custom spectral method solver and validated it against COMSOL MultiPhysics®. The response of the modeled sensor to a beam of particles is then simulated using Garfield++ and is compared to the data acquired in a beam test carried on in 2021 by the TimeSPOT Collaboration at the SPS, at CERN. Based on the results obtained with this simulation workflow, we conclude that reducing the resistivity of the graphitic columns remains the priority for significantly improving the time resolution of 3D diamond detectors. Once achieved, optimization of the detector geometry and readout electronics design will become equally important steps to further enhance the timing performance of these devices.
@@ 1 Introduction
The development of the ATLAS and CMS detectors for the High-Luminosity Large Hadron Collider [DaVia:2012ay, CMSTracker:2019aux, DallaBetta:2016czx, DallaBetta:2016dqn], the upcoming upgrade of the LHCb experiment [LHCB-TDR-023, LHCB-TDR-026], and the detectors at the hadronic version of the Future Circular Collider (FCC-hh) [FCC:2018vvp], or at the Muon Collider [InternationalMuonCollider:2024jyv] exemplify the growing demand for sensors capable of withstanding an unprecedented flux of ionizing particles.
@@
Similar challenges arise in technological and applied research, such as in the development of dosimeters for Flash Radiotherapy [Porter:2023crm, frontiers-review] or in providing precise process monitors in nuclear fission and fusion facilities [Metcalfe:2017slh].
@@
Physics requirements for experiments at future high luminosity colliders impose severe constraints especially on tracking detectors placed close to the interaction points. Besides very high radiation hardness, they must feature pixels with excellent time resolution to ensure an efficient tracking and micrometer-level precision in primary and secondary vertex reconstruction [LHCB-TDR-026].
@@
Among the many technologies that are being considered, like for example the recently proposed doping compensation in LGAD silicon sensors [Sola:2022dyc], diamond detectors have been extensively studied as highly promising candidates for addressing the challenges posed by extreme operating conditions. Diamond sensors offer indeed several advantages. Firstly, they have enhanced radiation tolerance due to the strong lattice binding energy and relatively low atomic number. Secondly, the high mobility and saturation velocity of charge carriers provide faster signal collection, leading to improved time resolution compared to silicon devices [Pomorski2009, pernegger]. Moreover, diamond sensors have a sufficiently large band-gap to make them solar-blind and operable at higher temperatures. This characteristic is particularly advantageous for applications where cooling systems are impractical due to space or complexity constraints.
@@
An intriguing feature of diamond is its ability to undergo laser-induced graphitization, transforming few of diamond around the laser focus into a conductive admixture of graphite and amorphous carbon. This process enables the fabrication of full-carbon sensors with high flexibility in the geometry of electrodes which can be realized either on the surface or in the bulk of the diamond crystals [Oliva:2019alx, PORTER2023109692, Anderlini:2021pei, Bachmair:2015iba].
@@
Similarly to the silicon sensors, the 3D geometry, where electrodes are perpendicular to the sensor surface, offers significant advantages over the traditional planar geometry. In this configuration, the drift length of charge carriers is substantially shorter with respect to the planar geometry, reducing the probability of charge trapping by radiation induced defects, and shortening the charge collection time, leading to faster signals. A sketch of a 3D electrode arrangement and a picture of a 3D diamond sensor are shown in Figure 2 and Figure 2 respectively.
@@
The fabrication of 3D diamond sensors presents substantial challenges due to the high refractive index of diamond, which degrades the laser focus due to spherical aberration, and because of the major resistance of the diamond bulk to laser-induced phase transitions. The transition from diamond to graphite within the bulk is complicated by differences in density between the two materials, leading to mechanical strain in the crystal lattice. Despite these challenges, significant progress has been made. For instance, Salter et al. demonstrated in 2017 that phase transitions in the diamond bulk occur through discrete nano-scale structural modifications aligned with the diagonals of the diamond lattice [salter2017high]. The electrical resistivity of these graphitized regions is influenced by random connections between conductive structures and is typically orders of magnitude higher than that of graphitized surfaces [Kononenko_2015].
@@
Several other studies, across various scientific communities, proposed methods to reduce the resistivity of the engraved structures. The most promising results have been achieved using femtosecond infrared lasers, where spherical aberration introduced by the diamond surface is corrected through Spatial Light Modulators [Anderlini:2022exf, test-beam-sensors]. This approach has enabled the fabrication of structures parallel to the diamond surface with a recorded resistivity of 0.02 [Salter:2014ohm]. Additionally, the same experimental setup has demonstrated the capability to engrave multiple electrodes simultaneously [Krueger:24]. On the other hand, employing Bessel femtosecond-pulsed beams without adaptive spherical aberration correction has resulted in the formation of fine electrodes orthogonal to the diamond surface, exhibiting a resistivity of 0.04 [KURIAKOSE2023110036].
@@
High resistivity in graphitized structures leads to considerable electrode resistance in diamond sensors. The increased resistance affects the signal propagation time through the electrodes. While the dependence of the propagation delay on the depth of the deposited charge can usually be neglected in the case of electrodes that can be assumed to be perfectly conductive (metallic electrodes hereafter), it becomes the dominant factor limiting the time resolution in 3D diamond sensors. Until recently, such a contribution to the time resolution from signal propagation was so large compared to those from carrier transport and electronic noise that could be effectively considered as the sole contribution, allowing for the use of simplified models. However, 3D laser-graphitized diamond sensors with a time resolution significantly below 100 ps have been reported [Anderlini:2022exf], thanks to the advanced graphitization techniques mentioned above, strongly motivating a renewed effort to model the signal formation in diamond detectors by properly taking into account the combined effect of carrier transport through the bulk and signal propagation through resistive electrodes.
@@
In this work, we propose an innovative numerical model based on an extension of the Ramo-Shockley theorem to conductive media to simulate the signal induction process in 3D diamond sensors with graphitic electrodes, including the contribution of the resistive elements as described in Refs. [Gatti1982, Riegler2002, Rigler:2e004jh, Riegler2019, Janssens:2890572].
@@
To validate our proposed method, we use experimental data obtained from a beam test conducted at CERN, thereby ensuring the reliability and applicability of our model. Having validated the method, we simulated sensors with different geometries and configurations to study their time resolution. The simulations allowed us to disentangle the various contributions to the time resolution, enabling a systematic investigation aimed at identifying potential optimizations.
@@ 2 Signal simulation in 3D diamond sensors with graphitic electrodes
Particles traversing a radiation detector release part of their energy inside the detection medium, generating charge carriers which drift towards the electrodes, inducing a current signal.
The simulation of signal generation involves several processes, which will be discussed in detail throughout this section:
1.
Interaction of an ionizing particle with the detector medium;
2.
Transport of the generated charge carriers through the medium;
3.
Propagation along the electrodes of the signal resulting from charge carrier motion;
4.
Amplification and shaping introduced by the readout front-end electronics.
This work does not introduce novelties in the simulation of the processes 1, 2 and 4, which are briefly reviewed in the following to provide a complete overview of simulation procedure and to report on the technical choices adopted in the deployment of the simulation. Process 3 is rarely relevant for sensors with metallic electrodes because the propagation of the signal through the perfectly conductive material is instantaneous compared to the time scale of process 2.
The simulation of process 1 and process 2 is presented in Section 2.1, while Section 2.2 reviews the description of the signal induction in the simplified case of metallic electrodes, by leveraging the Ramo-Shockley theorem.
As mentioned in the introduction, however, the high resistivity of graphitized electrodes makes the contribution to the time resolution from process 3 to dominate in the case of 3D diamond sensors.
To model the effect of process 3 on the induced signal we discuss two approaches:
approximating the effect of resistive electrodes with a discrete impedance network, following the equivalent circuit principle, hereafter referred to as (Section 2.3);
using a time-dependent extension of the Ramo-Shockley theorem (Section 2.4), in which two different methods to compute the electric field maps have been developed: a consolidated one based on the Finite Element Method (FEM, ), and an innovative one based on a pseudo-spectral method for solving the differential equations ().
In subsection 2.5, a comparison between all the simulation methods is provided, along with an explanation of the final steps related to noise and the simulation of the front-end electronics, required to process the output signals from all simulations so that they can be compared with experimental data for validation purposes.
Finally, to enable comparison of the simulation with experimental data, we discuss the simulation of the effects introduced by the electronics (process 4) in Section 2.6.
The simulation of processes 1 and 2 relies on the Garfield++ software package [Schindler:2012wta]. The simulation of process 3 with and have also been developed within the Garfield++ framework and were made available in the official repository of the project.
The sensor considered for the simulation is the 3D diamond sensor described in Ref. [Anderlini:2022exf], engraved in a 500 thick diamond specimen and featuring 12 -diameter graphitized electrodes, organized on a two-dimensional -pitch pixel matrix. The readout electrodes are arranged in square cells, with polarization electrodes placed at the four corners and the readout electrode at the center. Both polarization and readout electrodes are 450 long and are engraved starting from opposite sides of the crystal. Their resistance was indirectly measured to be 30 k, corresponding to a resistivity of 0.75 .
The sensor was selected for benchmarking the simulation based on a campaign of measurements conducted in 2022 using a 180 GeV pion beam, which provided experimental data for comparison with simulation results [Anderlini:2022exf].
@@ 2.1 Computing trajectories of charge carriers with Garfield++
The numerical model proposed in this work is built upon established methodologies for simulating the migration of charge carriers generated by the passage of ionizing particles. The motion of charges occurs within the electrostatic field defined by the polarization voltage applied to sensor electrodes.
The ionization charge generated by the traversing particle is simulated using the High Energy Electrodynamics (Heed) toolkit [Smirnov2005, Heed] as integrated in Garfield++. The detector response was investigated through simulations of tracks produced by 180 GeV/ pions impinging on the sensor surface and producing charge clouds of approximately 15000 electrons distributed in clusters along the particle path. The stochastic nature of the energy deposition is intrinsically taken into account in Heed, where the local fluctuations of the ionization charge along the particle track follow a Landau-like distribution [Landau:1944if]. This approach ensures a realistic representation of the spatial non-uniformity of the charge generation process within the sensor volume.
The drift of generated electrons and holes is determined by the electrostatic field, which determines both their trajectories toward the electrodes and their velocities, calculated by accounting for the carrier mobility () and saturation velocity () in diamond:
Due to the complex structure of the graphitized electrodes, carrier drift is neglected within the regions representing the conductive carbon phase. It should be noted that this simplification introduces discrepancies when the ionizing particle traverses an electrode because the deposited charge is not accounted for in our simulation model. However, beam test results on 3D diamond sensors show tracking efficiencies above 99%, indicating that the charge deposited in the graphitized material effectively contributes to the signals collected at the electrode ends [Anderlini:2022exf].
@@
The results described below were obtained computing the electrostatic potential in the diamond sensor with either a FEM implemented in COMSOL MultiPhysics® [comsol] on a custom mesh () [Janssens:2890572], or using an iterative spectral method on a uniform voxel grid, resetting the boundary conditions on voltage and charge density at each iteration ().
@@
2.2 Signal induced on ideal, perfectly conductive electrodes
According to the Ramo-Shockley theorem [Ramo1939, Shockley1938], the current induced on a given electrode by the drift of a charge along a path with velocity is given by
Here, denotes the weighting field, obtained by solving Laplace’s equation for the sensor geometry with a fixed potential applied to the electrode of interest, while all other electrodes are set to ground. Under static conditions, the electric potential is assumed to be uniform within each electrode, allowing the Dirichlet problem to be solved using the potential at the interface between the electrodes and the diamond as boundary conditions.
Neglecting the impedance of the electrodes, the weighting field can be considered as time-independent. In order to mark a distinction from the time-dependent weighting potential that will be discussed later, it is customary to refer to as the steady weighting field.
According to the prescription of the Ramo–Shockley theorem, this steady weighting field was evaluated using both numerical methods B and C. The resulting signal corresponds to the case in which the sensor electrodes can be considered metallic.
@@ 2.2 Signal induced on ideal, perfectly conductive electrodes
2.3 Modeling resistive electrodes with a discrete impedance network ()
As a preliminary step, in order to have a simplified benchmark available to gauge our results with, the system of a readout and a polarization graphitic electrode is described as a discrete impedance network, where the current signal induced by the drift of charge carriers inside the sensor is injected and propagated to the readout electronics input.
A crude approximation to take into account the effect of signal propagation through the graphitized electrodes is to model them as a discrete ladder of capacitors and resistors, as shown in Figure 5.
@@ 2.3 Modeling resistive electrodes with a discrete impedance network ()
From the equivalent circuit shown in Figure 5, a delta response function — i.e. the transfer function of the discrete-impedance network — is obtained by applying an elementary pulse in the current generators and measuring the current flowing towards ground. To account for the fluctuations of the ionization charge deposited through the depth of the columns (process 1), the elementary current pulse is split randomly through the discrete elements of the model according to the Landau distribution [Landau:1944if]. The transfer function is computed using the Ngspice software package [ngspice] and is then convolved numerically with the signal obtained with Garfield++ using the steady weighting field.
The discrete impedance network model provides a fast, heuristic but fairly reasonable estimate of the effect of the electrode resistivity on the signals [Anderlini:2021pei, Anderlini:2022exf]. However, it is limited to symmetric geometries and relies on idealized assumptions. First of all one has to consider the large uncertainties on the capacitance and the approximations coming in particular from neglecting the correlation effects between charge-carrier trajectories and signal propagation. In addition, the model is valid only insofar as the polarization and readout electrodes can be considered symmetric and that the shaping time constant of the readout electronics is of the same order of magnitude of the charge-carrier transport timescale, or larger.
In the following, we present a novel approach that, while computationally intensive and technically demanding, offers a more complete and theoretically robust insight into the physical processes underlying signal formation in 3D diamond sensors. Comparing predictions from this advanced method with those from the simplified model helps deepen the understanding of sensor behavior, especially in regimes where physical intuition still provides meaningful guidance.